Consider the following income statement:

Sales $558,400

Costs $346,800

Depreciation $94,500

EBIT ?

Taxes (35%) ?

Net Income ?

Fill in the missing numbers and then calculate the OCF. What is the depreciation tax shield?

2. An asset costs $545,000 and depreciated straight line to zero over eight years. The asset is to be used in a five-year project. At the end of the project the asset can be sold for $95,000. The tax rate is 35%. What is the after tax cash flow from the sale of this asset.

3. Cochran Inc. is considering a new three-year expansion project that requires an initial fixed asset investment of $1,950,000. The fixed asset will be depreciated straight-line to zero over its three-year tax life, after which time it will be worthless. The project is estimated to generate $2,145,000 in annual sales, which costs of $1,205,000. If the tax rate is 35%, what is the OCF for this project?

4. Calculate the NPV using the required return of 14% using the cash flows from the previous problem.

5. Do some sensitivity analysis. Suppose the president lowered the tax rate to 30%. Calculate the NPV again using a 14% required return.

Please help with solutions for 4. and 5.
Answer #1

1.

EBIT=558400-346800-94500=117100

Taxes=117100*35%=40985

net income=EBIT-taxes=117100-40985=76115

then calculate the OCF=EBIT+depreciation-taxes=117100+94500-40985=170615

the depreciation tax shield=depreciation*tax rate=94500*35%=33075